# The AI Imperative: Scaling High-Quality Peer Review in Machine Learning

### v1

@@ 4 The Vision: An AI-Augmented Peer Review Ecosystem
We envision a future where artificial intelligence (AI), particularly Large Language Models (LLMs), acts as an intelligent partner within the peer review ecosystem, collaborating with all human stakeholders to help mitigate the strains detailed in Section 2. This vision is not about replacing human judgment, which remains paramount, but about augmenting human capabilities. By automating or assisting with structured, repetitive, or cognitively demanding tasks, AI can free experts to concentrate on nuanced scientific assessment, ultimately enhancing the quality, efficiency, and fairness of peer review. This section first introduces key AI-driven tools and capabilities that form the foundation of this ecosystem. It then examines AI’s potential role from three key perspectives: assisting reviewers, authors, and area chairs (ACs), illustrating how these foundational tools are applied to empower each group.
@@ 6 Illustrative Experiments & Research Agenda
Position Supported: LLMs show promise in assisting with initial review tasks and predicting assessments, but In-Context Learning (ICL) with current data has limitations, highlighting needs for fine-tuning and richer, structured peer review data. Our experiments on ICLR corpora illustrate this.
Part 1: Review Component Generation. Task & Method: Given the initial submission, an LLM generated evaluation of the strengths/weaknesses of the paper. For rebuttals (given the reviews), the LLM identified key points to respond to. Compared against human data from ICLR 2024/25 OpenReview data using Recall (Avg Hits / Avg Real Points). Results & Interpretation (Table 1): LLMs showed higher recall for strengths (e.g., $0.927\pm 0.060$ for ICLR 2025) than weaknesses ($0.632\pm 0.000$), suggesting identifying flaws is harder and needs advanced alignment. Rebuttal point recall was high ($0.911\pm 0.040$). Increased recall in 2025 (see caption Table 1) may reflect LLM improvements and/or better parsing of increasingly AI-assisted review components (stylistically similar to output from advanced LLMs like OpenAI’s models). This highlights LLM utility and the evolving, AI-infused data landscape, stressing need for human oversight for nuanced evaluation.
Part 2: Rating Prediction. Task & Method: Using ICL on ICLR 2025 data with varying examples ($n=0,1,2,3$), an LLM predicted initial/final ratings and score changes. Evaluated by MAE/RMSE. Results & Interpretation (Table 2): LLMs offer a reasonable baseline (e.g., initial rating MAE $2.2857\pm 0.0095$ with $n=2$; final rating MAE $0.6709\pm 0.0052$ with $n=1$). Scaling $n$ shows diminishing returns, implying ICL limits for complex regression without explicit training on underlying factors. Significant gains likely require fine-tuning on larger, specialized datasets with structured rationale. Inherent subjectivity in peer review may also cap accuracy. Score change prediction was particularly hard. LLMs offer estimations, but precise prediction needs advancement and richer data; human judgment remains indispensable. These experiments underscore LLM assistive potential but also highlight limitations, reinforcing calls for human oversight, sophisticated AI development (including fine-tuning on richer data), and deeper understanding of review process data.
Table 1: LLM recall in extracting key points from ICLR peer-review corpora (2024 vs. 2025). Values are mean $\pm$ 95% CI. Increased recall for ICLR 2025 may reflect evaluating LLM improvements and/or better parsing of increasingly common AI-assisted review components (e.g., content stylistically similar to that from advanced LLMs such as OpenAI’s models).
@@
Table 2: Effect of in-context examples ($n$) on LLM rating prediction accuracy (ICLR 2025 data). MAE/RMSE (mean $\pm$ 95% CI); lower is better.
@@ 8 Discussion and Conclusion
The ML community faces a critical juncture: submission volumes threaten traditional peer review. This paper posits that AI assistance is an imperative. We envision LLMs as collaborators augmenting authors, reviewers, and ACs, helping manage burdens, enhance rigor, promote consistency, guide development, and support decisions. Our illustrative experiments show nascent LLM potential. Transforming this potential into a robust ecosystem critically depends on access to richer, more granular, ethically-sourced data on the peer review process itself—especially on deliberation dynamics and the reasoning behind evaluative judgments. Current datasets lack this depth. This requires concerted community action: (1) Conference organizers and platforms (e.g., OpenReview) must prioritize structured collection of detailed deliberation data, with privacy safeguards and minimal burden. (2) Researchers must develop AI tools for augmentation, keeping humans in control. (3) The community must proactively address ethical challenges (bias, misuse, privacy) with clear guidelines. The path forward involves pilot programs, transparent development, community engagement for trust, and continuous evaluation. Empowering human experts, not diminishing them, is the goal. Investing in AI assistance and supporting data infrastructure can build a more scalable, robust, and fair peer review system, safeguarding ML research integrity and progress.
@@ 9 Limitations and Future Work
This position paper acknowledges limitations. Our illustrative experiments are conceptual and need rigorous, large-scale empirical validation. Key limitations: (1) Empirical Grounding and Scalability: Proposed AI assistants require substantial research, data engineering, and validation beyond our illustrations. (2) Scientific Reasoning Complexity: Current LLMs struggle with deep scientific reasoning, novelty assessment, and understanding complex, implicit argumentation. (3) Data Availability, Quality, and Heterogeneity: Richer, structured data are not yet widely available; integration and quality assurance are major hurdles. (4) Unintended Consequences: Introducing AI into peer review can have unforeseen effects (e.g., over-reliance, de-skilling, new gaming strategies). Future work must develop unified, privacy-preserving data schemas for fine-tuning specialized, verifiable AI models capable of claim checking and decision support with clear rationales. This necessitates intuitive human-AI collaboration interfaces and continuous ethical auditing (including bias detection and fairness metrics) to ensure trust. Longitudinal trials across venues are vital to measure AI’s impact on review quality, efficiency, and diversity, while advancing robust content provenance and misuse mitigation (e.g., watermarking, resilient detectors) is key to safeguarding legitimate AI assistance.
Acknowledgements We thank Hao Sun for helpful brainstorming discussions and advice provided throughout the project. This work was supported by Azure sponsorship credits from Microsoft’s AI for Good Research Lab and Microsoft’s Accelerate Foundation Models Academic Research Initiative. QW is funded by GSK, and SH is funded by AstraZeneca. We also acknowledge data collected from Paper Copilot, an independent initiative funded by Jing Yang.
@@ B.1 General Considerations for ICL
For all In-Context Learning (ICL) experiments, we would conceptually use publicly available data from OpenReview for conferences like ICLR (e.g., ICLR 2024 & 2025 data, focusing on papers with full review cycles). Experiments are run with state-of-the-art LLMs available at the time of execution (e.g., O3 from OpenAI)

### v2

@@ 4 The Vision: An AI-Augmented Peer Review Ecosystem
We envision a future where artificial intelligence (AI), particularly LLMs, acts as an intelligent partner within the peer review ecosystem, collaborating with all human stakeholders to help mitigate the strains detailed in Section 2. This vision is not about replacing human judgment, which remains paramount, but about augmenting human capabilities. By automating or assisting with structured, repetitive, or cognitively demanding tasks, AI can free experts to concentrate on nuanced scientific assessment, ultimately enhancing the quality, efficiency, and fairness of peer review. This section first introduces key AI-driven tools and capabilities that form the foundation of this ecosystem. It then examines AI’s potential role from three key perspectives: assisting reviewers, authors, and area chairs (ACs), illustrating how these foundational tools are applied to empower each group.
@@ 6 Illustrative Experiments & Research Agenda
Position Supported: LLMs show promise in assisting with initial review tasks and predicting assessments, but In-Context Learning (ICL) with current data has limitations, highlighting needs for fine-tuning and richer, structured peer review data. Our experiments on ICLR corpora illustrate this—we provide experimental implementation details in Appendix B.
Part 1: Review Component Generation. Task & Method: Using few-shot prompting, an LLM was tasked to generate strengths and weaknesses from a paper’s content, and separately, to identify key rebuttal points from initial reviews. We evaluated the semantic overlap of the LLM’s output against the actual human-written content from ICLR 2024/25 OpenReview data, measuring recall via an LLM-as-judge protocol (Avg Hits / Avg Real Points). Results & Interpretation (Table 1): LLMs showed higher recall for strengths (e.g., $0.927\pm 0.060$ for ICLR 2025) than weaknesses ($0.632\pm 0.000$), suggesting identifying flaws is harder and needs advanced alignment. Rebuttal point recall was high ($0.911\pm 0.040$). Increased recall in 2025 (see caption Table 1) may reflect LLM improvements and/or better parsing of increasingly AI-assisted review components (stylistically similar to output from advanced LLMs). This highlights LLM utility and the evolving, AI-infused data landscape, stressing need for human oversight for nuanced evaluation.
Part 2: Rating Prediction. Task & Method: Using few-shot In-Context Learning (ICL) with a varying number of examples ($n=0,1,2,3$), an LLM predicted initial ratings (from paper content), final ratings (from reviews and author rebuttals), and the resulting score change. We prompted the model with text from the ICLR 2025 dataset and evaluated its predictions against the ground-truth scores using Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE). Results & Interpretation (Table 2): LLMs offer a reasonable baseline (e.g., initial rating MAE $2.2857\pm 0.0095$ with $n=2$; final rating MAE $0.6709\pm 0.0052$ with $n=1$). Scaling $n$ shows diminishing returns, implying ICL limits for complex regression without explicit training on underlying factors. Significant gains likely require fine-tuning on larger, specialized datasets with structured rationale. Inherent subjectivity in peer review may also cap accuracy. Score change prediction was particularly hard. LLMs offer estimations, but precise prediction needs advancement and richer data; human judgment remains indispensable. These experiments underscore LLM assistive potential but also highlight limitations, reinforcing calls for human oversight, sophisticated AI development (including fine-tuning on richer data), and deeper understanding of review process data.
Table 1: LLM recall in extracting key points from ICLR peer-review corpora (2024 vs. 2025). Values are mean $\pm$ 95% Confidence Intervals (CI). Increased recall for ICLR 2025 may reflect evaluating LLM improvements and/or better parsing of increasingly common AI-assisted review components (e.g., content stylistically similar to that from advanced LLMs).
@@
Table 2: Effect of in-context examples ($n$) on LLM rating prediction accuracy (ICLR 2025 data). Mean Absolute Error (MAE) / Root Mean Squared Error (RMSE) (mean $\pm$ 95% CI); lower is better.
@@ 8 Discussion and Conclusion
The ML community faces a critical juncture: submission volumes threaten traditional peer review. This paper posits that AI assistance is an imperative. We envision LLMs as collaborators augmenting authors, reviewers, and ACs, helping manage burdens, enhance rigor, promote consistency, guide development, and support decisions. Our illustrative experiments show nascent LLM potential. Transforming this potential into a robust ecosystem critically depends on access to richer, more granular, ethically-sourced data on the peer review process itself—especially on deliberation dynamics and the reasoning behind evaluative judgments. Current datasets lack this depth. This requires concerted community action: (1) Conference organizers and platforms must prioritize the structured collection of detailed deliberation data. Platforms such as OpenReview, for example, provide an invaluable foundation through their public-facing nature and data APIs, but could be enhanced to capture more granular interactions (e.g., explicitly linking score changes to specific rebuttal arguments), all while maintaining robust privacy safeguards and minimizing user burden. (2) Researchers must develop AI tools for augmentation, keeping humans in control. (3) The community must proactively address ethical challenges (bias, misuse, privacy) with clear guidelines. The path forward involves pilot programs, transparent development, community engagement for trust, and continuous evaluation. Empowering human experts, not diminishing them, is the goal. Investing in AI assistance and supporting data infrastructure can build a more scalable, robust, and fair peer review system, safeguarding ML research integrity and progress.
@@ 9 Limitations and Future Work
This position paper acknowledges limitations. Our illustrative experiments are conceptual and need rigorous, large-scale empirical validation. Key limitations: (1) Empirical Grounding and Scalability: Proposed AI assistants require substantial research, data engineering, and validation beyond our illustrations. (2) Scientific Reasoning Complexity: Current LLMs struggle with deep scientific reasoning, novelty assessment, and understanding complex, implicit argumentation. (3) Data Availability, Quality, and Heterogeneity: Richer, structured data are not yet widely available; integration and quality assurance are major hurdles. While platforms like OpenReview have democratized access to review data, the information often remains in semi-structured free-text, limiting the development of AI that can deeply reason about the deliberative process. (4) Unintended Consequences: Introducing AI into peer review can have unforeseen effects (e.g., over-reliance, de-skilling, new gaming strategies). Future work must develop unified, privacy-preserving data schemas for fine-tuning specialized, verifiable AI models capable of claim checking and decision support with clear rationales. For instance, this could involve collaboration with platforms like OpenReview to pilot new review interfaces that prompt for structured rationale for score changes or allow for fine-grained annotation of claims within discussion threads. This necessitates intuitive human-AI collaboration interfaces and continuous ethical auditing (including bias detection and fairness metrics) to ensure trust. Longitudinal trials across venues are vital to measure AI’s impact on review quality, efficiency, and diversity, while advancing robust content provenance and misuse mitigation (e.g., watermarking, resilient detectors) is key to safeguarding legitimate AI assistance.
Acknowledgements We thank Andrew McCallum for insightful discussions, valuable feedback, and high-level guidance that significantly enhanced the quality and impact of this work. We are also grateful to Hao Sun for helpful brainstorming discussions and advice provided throughout the project. This research was supported by Azure sponsorship credits provided by Microsoft’s AI for Good Research Lab and the Accelerate Foundation Models Academic Research Initiative. QW is supported by funding from GSK, and SH is supported by AstraZeneca. Additionally, we acknowledge the data collected from Paper Copilot, an independent initiative funded by Jing Yang.
@@ B.1 General Considerations for ICL
For all In-Context Learning (ICL) experiments, we would conceptually use publicly available data from OpenReview for conferences like ICLR (e.g., ICLR 2024 & 2025 data, focusing on papers with full review cycles). Experiments are run with a state-of-the-art LLM available at the time of execution; specifically, we used OpenAI’s O3 model throughout.